OpenStudy (anonymous):

Help Please!!!!!!!! Ill Give Medal!!!!! Solve. 4x - 6 = -3x + 8 x = ??????

OpenStudy (anonymous):

4x - 6 = -3x + 8

OpenStudy (shaik0124):

4x-6-4x=-3x-4x+8

OpenStudy (anonymous):

4x+3x=8+6 7x=14 x=14/7

OpenStudy (anonymous):

:D thanks :D

OpenStudy (shaik0124):

-6=-7x+8 -6-8=-7x+8-8 -14=-7x x=14/7

OpenStudy (anonymous):

Josh has been running for 6 min and completes a lap in 7 min. Gabe has been running for 2 min and completes a lap in 9 min. How many more laps must Josh and Gabe complete to have run the same amount of time? ?????? ( not multiple choice ) ( i just don't feel like doing the math right now )

OpenStudy (anonymous):

??? are you still there ???

OpenStudy (anonymous):